# Online Multivariate Regularized Distributional Regression for High-dimensional Probabilistic Electricity Price Forecasting

### v1

@@ 4 Results
Figure 8: Illustrative simulations drawn from two models.
Scores
Table 4 gives the results for the scoring rules for each model and Figure 9 provides one-side Diebold-Mariano-tests for all pairwise model comparisons. As a mental guidance for the increasing complexity of the online regression models, remember Section 3: The first two models have an adaptive, but unconditional estimation for the scale parameter/matrix under a Gaussian assumption. The Copula-based models employ online, conditional estimation for all marginal distributional parameters and an adaptive, but unconditional estimation for the dependence structure, while the multivariate distributional regression models yield an online estimation conditional multivariate distribution. Let us note a few main results from Table 4 here:
  • The baseline LARX-models yield the best performance in terms of the RMSE. This is common theme in probabilistic forecasting using distributional models due to the fact that the likelihood-based estimation down-weights observations with high (estimated) variance, thereby reducing the precision in the mean estimation to improve the distributional fit overall. Similar results can be observed e.g. in Marcjasz et al. (2023); Hirsch et al. (2024). The LARX model with the multivariate Gaussian distribution also yields, across all various metrics, very robust results.
  • The two univariate models, the $\operatorname{LARX}+\mathcal{N}(0,\sigma)$ and the univariate distributional regression model without Copula yield, not surprisingly, weak scores for the multivariate scoring rules (VS, ES, DSS and LS), while being competitive in the univariate scores with their respective counterparts. This result underscores that univariate models cannot capture the dependence structure and therefore miss a crucial element for probabilistic forecasting of electricity prices.
  • The univariate online distributional regression including the copula yields strong results for VS and the ES. However, for the DSS and the LS, the Copula-based models do not perform as well. An issue here might be necessary two-step approach, which naturally introduces some friction in the estimation of the scale matrix if the marginal assumption does not fit perfectly.
  • For the multivariate online distributional regression models, we see that the models using the Cholesky-based parametrization provide better performance. This is likely due to the fact that the regularized CD is closer to the natural, time-based structure of the (conditional) covariance than the LRA. This is also visible in Figure 8, which visually compares trajectories from the two scale matrix parameterizations.
  • Overall, the estimation using LASSO compared to OLS increases the forecasting performance for the multivariate distributional regression models significantly. The comparison of the independence configuration of the distributional regression models and their unrestricted counterparts show that the CD-based models yield higher RMSE errors, most likely through cross-transmission of errors in the covariance matrix.
  • The $p$-values of the DM-test in Figure 9 largely confirm the statistical significance of the aforementioned results. We note the strong performance of the Copula-based models for the ES and the statistically significant superior performance of the multivariate distributional regression for the DSS and LS.
Overall, our results highlight that neglecting the dependence structure by relying solely on marginal, univariate models yields subpar probabilistic forecasting performance. We note that for the truly multivariate approaches, using both Copula-based combinations of univariate models and the fully multivariate distributional regression yield statistically significant performance improvements. However, the CRPS and Energy Score are not improved by the multivariate models. This might be due to the fact that the dependence structure in the day-ahead electricity prices does not change significantly over time, as well as error transmission in modelling the conditional scale matrix.
Model RMSE MAE CRPS $\text{VS}_{p=0.5}$ $\text{VS}_{p=1}$ ES DSS LS
LARX + N(0, $\sigma$) 7.346 4.530 3.406 1.017 6.373 20.935 119.170 81.510
LARX + N(0, $\Sigma$) 7.346 4.532 3.406 0.885 5.712 20.455 84.487 64.149
oDistReg 7.463 4.393 3.305 1.012 6.546 20.701 114.483 76.754
oDistReg+GC 7.454 4.394 3.323 0.851 5.548 20.252 103.057 61.536
oDistReg+spGC 7.449 4.391 3.320 0.851 5.548 20.238 102.207 61.476
oMvDistReg(t, CD, OLS, ind) 8.027 4.447 3.456 0.851 5.495 21.351 137.906 70.410
oMvDistReg(t, LRA, OLS, ind) 8.014 4.440 3.440 0.855 5.518 21.276 134.802 70.083
oMvDistReg(t, CD, OLS) 8.345 4.723 3.559 0.861 5.574 21.762 89.068 55.861
oMvDistReg(t, LRA, OLS) 8.165 4.605 3.556 0.881 5.656 21.980 135.251 70.280
oMvDistReg(t, CD, LASSO) 8.323 4.664 3.494 0.870 5.637 21.420 81.496 55.356
oMvDistReg(t, LRA, LASSO) 8.858 5.190 3.929 0.992 6.433 24.045 128.334 73.295
@@
Table 4: Scoring Rules for the full out of sample period of 736 days. The best score in each column is marked bold. Note that the $\operatorname{LARX}+\mathcal{N}(0,\sigma)$, the oDistReg and the oMvDistreg(…, ind) models do not model the dependence structure.
@@
Computation times
Table 5 gives computation times for all experiments. The initial fit for the multivariate distributional regression model takes a few minutes, the update algorithm can be executed in seconds. On a standard laptop, the experiments can be run in about 2 hours. Even though we did not run a repeated batch fitting, an estimate for the benefit of online vs. repeated batch fitting can be achieved by multiplying the initial fit duration with 736 days of out of sample and comparing this to the total time of the online study:
Speedup=Initial $\text{Speedup}=\frac{\text{Initial Fit}\times T}{\text{Total Time}}.$
By this (albeit simple) measure, the online learning improves computation by a factor of 80 to 400. Thereby our online update algorithm makes the approach practically viable for researchers and data scientists without access to specialized high-performance computation centers. These estimates are in line with benefits reported in Hirsch et al. (2024) for the univariate online distributional regression case in an explicit comparison.
Model Initial Fit Avg. Update Std. Update Total Time Est. Speedup
LARX + N(0, $\sigma$) 1.74 0.02 0.04 13.70 $\times$93
LARX + N(0, $\Sigma$) 1.74 0.02 0.04 13.70 $\times$93
oDistReg 25.08 0.12 0.07 114.26 $\times$161
oDistReg+GC 25.08 0.12 0.07 114.71 $\times$160
oDistReg+spGC 25.28 0.29 0.07 241.68 $\times$76
oMvDistReg(t, CD, OLS, ind) 46.44 0.09 0.07 113.68 $\times$300
oMvDistReg(t, LRA, OLS, ind) 55.32 0.11 0.03 132.70 $\times$306
oMvDistReg(t, CD, OLS) 146.94 0.31 0.01 377.33 $\times$286
oMvDistReg(t, LRA, OLS) 188.24 1.96 0.17 1628.10 $\times$85
oMvDistReg(t, CD, LASSO) 290.06 3.07 0.03 2549.02 $\times$83
oMvDistReg(t, LRA, LASSO) 864.99 0.68 0.04 1363.32 $\times$466

### v2

@@ 4 Results
4.1 Forecasting the Time-varying Dependence Structure
Seven exemplary predicted covariance matrices are shown in Figure 10 and some time-varying behavior of the covariance matrix over the week, especially in the morning hours is visible. This section aims to further analyze the time-varying dependence structure. Figure 11 plots the evolution of the predicted volatility (standard deviation) and correlation matrix for the oMvDistReg($t$, MCD ,OLS) model over time. The lower three panels show the 1st, 2nd and 3rd off-diagonal of the correlation matrix over time. We see that the correlation is the lowest around the hours 5-7, which corresponds to start of the morning ramp. The correlation between to hours $h$ and $h+i$ decays stronger for days with lower levels of volatility and vice versa. There are some weekly patterns visible: The dependence between the hours 5-7 and 16-19 tapers off more strongly on working days. This is likely driven by stronger shapes in prices and residual demand. Overall, we see that the dependence structure is not constant over time, which underlines the importance of including a time-varying dependence structure in multivariate electricity price forecasting.
Figure 11: Evolution of the predicted volatility (standard deviation) and correlation matrix for the oMvDistReg($t$,MCD,OLS) model over time. The lower three panels show the 1st, 2nd and 3rd off-diagonal of the correlation matrix over time. We see that the correlation is the lowest around the hours 5-6, which corresponds to start of the morning ramp. The correlation between to hours $h$ and $h+i$ decays stronger for days with lower levels of volatilty and vice versa. There are some weekly patterns visible.
4.2 Forecasting Accuracy
Panel A: Full sample (2018-12-27 to 2023-12-31) RMSE MAE CRPS $\text{VS}_{p=0.5}$ $\text{VS}_{p=1}$ ES DSS LS $\text{MC}_{0.95}$ $\text{JPBW}_{0.95}$ LARX+N(0, $\sigma$) 31.177 18.214 13.677 4.709 887.202 81.201 176.367 110.038 -0.134 131.129 LARX+N(0, $\Sigma$) 31.182 18.219 13.678 3.396 654.953 78.603 124.345 84.057 -0.161 118.914 LARX+Adaptive CP 31.268 18.374 14.402 -0.655 53.751 LARX+GARCH(1,1) 31.181 18.218 12.992 4.536 850.636 78.312 159.545 101.790 -0.040 140.459 oDistReg 36.040 19.104 13.980 4.722 1046.549 84.445 160.422 99.266 0.015 332.441 oDistReg+GC 33.901 18.511 13.501 3.591 793.156 78.372 124.869 77.934 -0.001 294.100 oDistReg+spGC 33.964 18.522 13.508 3.597 793.049 78.417 124.646 77.964 0.004 294.163 oMvDistReg(t, CD, OLS, ind) 38.898 20.207 15.410 3.787 771.899 89.011 174.092 95.446 -0.046 191.574 oMvDistReg(t, CD, OLS) 46.742 23.493 17.463 3.747 797.964 97.948 126.357 76.013 -0.017 273.946 oMvDistReg(t, CD, LASSO) 46.793 23.474 17.612 3.938 872.181 98.819 127.551 75.598 0.002 306.158 oMvDistReg(t, MCD, OLS, ind) 38.937 20.196 15.379 3.861 792.223 88.771 175.635 95.274 -0.034 196.792 oMvDistReg(t, MCD, OLS) 46.450 23.376 17.255 3.596 747.954 96.739 121.365 74.907 -0.037 234.422 oMvDistReg(t, MCD, LASSO) 46.581 23.345 17.261 3.699 787.099 96.803 124.217 74.789 0.017 298.814 oMvDistReg(t, LRA, OLS, ind) 39.197 20.292 15.710 5.189 >10,000>10,000 93.714 175.884 95.602 -0.049 201.897 oMvDistReg(t, LRA, OLS) 41.642 21.292 16.520 3.940 1161.734 94.398 174.901 93.993 -0.072 176.183 oMvDistReg(t, LRA, LASSO) 41.725 21.328 16.419 3.911 963.429 93.844 169.019 93.780 -0.047 188.270
Panel B: First sub-sample (2018-12-27 to 2020-12-31, $n=736$) RMSE MAE CRPS $\text{VS}_{p=0.5}$ $\text{VS}_{p=1}$ ES DSS LS $\text{MC}_{0.95}$ $\text{JPBW}_{0.95}$ LARX+N(0, $\sigma$) 7.568 4.790 3.588 1.163 59.095 21.940 120.417 82.117 -0.063 39.127 LARX+N(0, $\Sigma$) 7.568 4.791 3.588 0.912 52.607 21.368 85.142 64.467 -0.080 36.139 LARX+Adaptive CP 7.751 4.959 3.831 -0.525 18.469 LARX+GARCH(1,1) 7.567 4.791 3.601 1.258 65.391 22.099 118.522 81.369 -0.049 42.002 oDistReg 7.849 4.552 3.399 0.994 56.194 21.024 112.649 74.195 -0.006 50.476 oDistReg+GC 7.828 4.551 3.394 0.877 53.687 20.579 76.162 56.599 -0.025 43.055 oDistReg+spGC 7.819 4.548 3.390 0.877 53.706 20.557 75.768 56.561 -0.021 43.102 oMvDistReg(t, CD, OLS, ind) 8.034 4.664 3.557 0.952 56.556 21.777 133.695 71.549 -0.061 44.810 oMvDistReg(t, CD, OLS) 8.050 4.713 3.536 0.916 55.459 21.373 84.530 56.039 -0.045 49.244 oMvDistReg(t, CD, LASSO) 8.050 4.671 3.525 0.927 56.211 21.346 84.322 55.780 -0.012 56.089 oMvDistReg(t, MCD, OLS, ind) 8.056 4.659 3.509 0.961 56.386 21.557 122.110 70.944 -0.017 49.627 oMvDistReg(t, MCD, OLS) 8.059 4.699 3.505 0.900 54.724 21.277 80.581 55.291 -0.036 45.742 oMvDistReg(t, MCD, LASSO) 8.080 4.674 3.500 0.920 55.525 21.281 81.532 55.158 0.007 58.907 oMvDistReg(t, LRA, OLS, ind) 8.047 4.662 3.564 0.941 56.963 21.830 126.458 71.156 -0.041 44.971 oMvDistReg(t, LRA, OLS) 8.057 4.681 3.566 0.952 56.712 21.840 120.469 69.491 -0.034 45.885 oMvDistReg(t, LRA, LASSO) 8.079 4.671 3.552 0.959 57.035 21.798 120.635 69.375 -0.021 48.611
Panel C: Second sub-sample (2021-01-01 to 2023-12-31, $n=1095$) RMSE MAE CRPS $\text{VS}_{p=0.5}$ $\text{VS}_{p=1}$ ES DSS LS $\text{MC}_{0.95}$ $\text{JPBW}_{0.95}$ LARX+N(0, $\sigma$) 39.835 27.237 20.459 7.093 1443.811 121.033 213.973 128.804 -0.182 192.967 LARX+N(0, $\Sigma$) 39.841 27.245 20.460 5.065 1059.817 117.073 150.695 97.225 -0.215 174.550 LARX+Adaptive CP 39.931 27.391 21.507 -0.742 77.465 LARX+GARCH(1,1) 39.840 27.243 19.304 6.739 1378.435 116.095 187.118 115.516 -0.034 206.636 oDistReg 46.157 28.885 21.092 7.228 1712.213 127.073 192.533 116.117 0.029 521.962 oDistReg+GC 43.365 27.894 20.295 5.415 1290.187 117.218 157.607 92.275 0.014 462.839 oDistReg+spGC 43.450 27.915 20.309 5.425 1289.995 117.308 157.499 92.349 0.021 462.912 oMvDistReg(t, CD, OLS, ind) 49.867 30.654 23.377 5.693 1252.714 134.201 201.244 111.508 -0.035 290.222 oMvDistReg(t, CD, OLS) 60.081 36.115 26.825 5.650 1297.036 149.418 154.471 89.439 0.002 424.979 oMvDistReg(t, CD, LASSO) 60.148 36.111 27.081 5.962 1420.631 150.892 156.608 88.918 0.012 474.241 oMvDistReg(t, MCD, OLS, ind) 49.915 30.639 23.357 5.810 1286.812 133.948 211.612 111.628 -0.045 295.708 oMvDistReg(t, MCD, OLS) 59.700 35.929 26.497 5.407 1213.906 147.461 148.778 88.092 -0.038 361.243 oMvDistReg(t, MCD, LASSO) 59.869 35.895 26.510 5.567 1278.824 147.565 152.907 87.985 0.024 460.066 oMvDistReg(t, LRA, OLS, ind) 50.255 30.798 23.874 8.044 >10,000>10,000 142.030 209.106 112.033 -0.054 307.374 oMvDistReg(t, LRA, OLS) 53.441 32.456 25.226 5.949 1904.470 143.167 211.488 110.463 -0.098 263.763 oMvDistReg(t, LRA, LASSO) 53.547 32.524 25.068 5.894 1572.657 142.270 201.539 110.184 -0.064 282.142
@@ 4.2 Forecasting Accuracy
Table 4: Scoring Rules. The best score in each column is marked bold. Note that the $\operatorname{LARX}+\mathcal{N}(0,\sigma)$, the oDistReg and the oMvDistreg(…, ind) models do not model the dependence structure.
Table 4 gives the results for the scoring rules for each model. Figure 12 gives daily skills scores of the multivariate distributional regression models over the Copula-based benchmarks. Figure 13 provides one-sided Diebold-Mariano-tests for all pairwise model comparisons. As a mental guidance for the increasing complexity of the online regression models, remember Section 3: The first three models have an adaptive, but unconditional estimation for the scale parameter/matrix under a Gaussian assumption. The Copula-based models employ online, conditional estimation for all marginal distributional parameters and an adaptive, but unconditional estimation for the dependence structure, while the multivariate distributional regression models yield an online estimated conditional multivariate distribution.
Let us note a few main results here, before we discuss the results in more detail.
  • 1.
    Calibration: In terms of the calibration of the joint prediction bands, the multivariate distributional regression models and the Copula-based models yield the best performance. The univariate models yield poor performance with up to 10-16% under-coverage the 95%-JPB for the Gaussian-based models and, given we can only use marginal, quantile-based prediction bands for conformal prediction, 65% under-coverage for the LARX-ACP model. For the models yielding calibrated JPBs, the multivariate distributional regression models yield the narrowest JPBs.
  • 2.
    Marginal Scores: Here, the LARX models yield the best performance in terms of RMSE, MAE and CRPS. These models yield sharp predictions, but at the cost of poor coverage of the joint prediction bands. This is also visible in the very small JPBW.
  • 3.
    Multivariate Scores: The Copula-based models yield the best performance in terms of the ES. The MCD-based multivariate distributional regression model yields the best overall performance in terms of the LS and DSS as well as a second-best performance in the VS.
The $p$-values of the DM-test in Figure 13 largely confirm the statistical significance of the aforementioned results. We note the strong performance of the Copula-based models for the ES and the statistically significant superior performance of the multivariate distributional regression for the DSS, LS, and the VS. We note the disconnect between the results of the ES on one hand and the VS, DSS and LS on the other hand, which is in line with the findings of Pinson and Tastu (2013); Marcotte et al. (2023) and Alexander et al. (2024).
Looking at Figure 12, we see that skill scores are also well correlated with the current market regime. Skill scores are defined as
$\operatorname{SS}_{\text{model}}=1-\frac{\operatorname{LS}_{\text{model}}}{\operatorname{LS}_{\text{baseline}}},$
where we use the oDist+GC model as baseline. During the period of high and volatile prices in 2022, the skill scores of the multivariate distributional regression models over the Copula-based benchmarks are at their highest, while during the low-price period in 2020, the skill scores are at their lowest. Additionally, we see that the models assuming independence have have an increasingly better performance in recent years. This can be interpreted as a sign that the dependence structure has both weakened and changed with the changing market conditions.
Figure 12: Rolling 182-day average of skill scores. The skill score is defined as $\operatorname{SS}=1-\operatorname{LS}_{\text{model}}/\operatorname{LS}_{\text{baseline}}$, where the baseline is the oDist+GC model. A positive skill score indicates an improvement over the baseline. The secondary $y$-axis shows the daily average day-ahead price.
A common theme for distributional regression models is the trade-off between accuracy in the point predictions and the distributional error metrics. Marcjasz et al. (2023) and Hirsch et al. (2024) note this observation in the univariate case, trading RMSE for CRPS. This behavior is rooted in the fact that observations with predicted high variance are weighted down in the estimation of the mean parameters, which leads to slightly worse point predictions, but better distributional forecasts. At the same time, these observations with high variance are especially costly in terms of the RMSE. We see that the behaviour extends in the multivariate case as well, trading off marginal accuracy somewhat better modeling of the dependence structure in otherwise equal models. So far, this issue has not been discussed for multivariate distributional regressions and yet deserves further attention in future research (Gioia et al., 2022; Muschinski et al., 2022).
For the multivariate online distributional regression models, we see that the models using the Cholesky-based parametrizations provide better performance. This is likely due to the fact that these parametrizations are closer to the natural, time-based structure of the (conditional) covariance than the LRA and therefore the path-based regularization allows to to select a parsimounous model, which still reflects the shape of the dependence structure. We further discuss the regularization in the following Subsection 4.3. On the other hand, the estimation using LASSO yields slight improvements in the scoring rules. The rather marginal gain might be explained by the fact that all fundamental variables are known to be relevant for electricity price formation, and hence the regularization does not remove many variables. However, the regularization might help to stabilize the estimates in the online learning setting.
@@
Additionally, we analyze the performance of the models for extreme price spikes and for days with large spreads during the day. These days are interesting from a risk-management respectively from a battery optimization perspective. We define a price spike as a day with a minimum or maximum price exceeding the 5% resp. 95%-quantile of all minimum resp. maximum prices. A large spread event is defined as the min-to-max spread exceeding the 90%-quantile of all min-to-max spreads. Formally:
$\operatorname{Spike}$ $=\{t\mid\min(\bm{\mathrm{y}}_{t})<\operatorname{Q}_{0.05}(\min(\bm{\mathrm{y}}_{t}))\lor\max(\bm{\mathrm{y}}_{t})>\operatorname{Q}_{0.05}(\max(\bm{\mathrm{y}}_{t}))\},\quad\text{and}$
$\operatorname{Spread}$ $=\{t\mid\max(\bm{\mathrm{y}}_{t})-\min(\bm{\mathrm{y}}_{t})>\operatorname{Q}_{0.90}(\max(\bm{\mathrm{y}}_{t})-\min(\bm{\mathrm{y}}_{t}))\}.$
The results for the CRPS, LS and the miscoverage are given in Table 5. For the CRPS, we see that the ordering of the models is similar to the overall results, and remains largely unchanged between spike and no-spike events. For the miscoverage, we see that the multivariate distributional regression models yield the best results for both regimes for large spread events, and the best results for the price spike events, while the Copula-based models yield the best results for the no-spike events. Overall, these results highlight the robustness of the distributional regression models also for extreme events.
CRPS LS $\text{MC}_{0.95}$
Spread Spike Spread Spike Spread Spike
Low High Low High Low High Low High Low High Low High
LARX+N(0, $\sigma$) 3.53 (05) 37.85 (02) 10.93 (03) 29.02 (03) 80.43 (15) 169.63 (15) 100.06 (15) 157.46 (15) -0.06 (13) -0.37 (14) -0.10 (14) -0.45 (13)
LARX+N(0, $\Sigma$) 3.53 (06) 37.85 (03) 10.92 (02) 29.02 (02) 63.70 (07) 123.39 (07) 77.75 (07) 119.74 (07) -0.06 (15) -0.43 (15) -0.12 (15) -0.52 (15)
LARX+Adaptive CP 3.67 (16) 38.27 (04) 11.25 (06) 30.16 (04) (17) (17) (17) (17) -0.56 (16) -0.79 (16) -0.56 (16) -0.86 (16)
LARX+GARCH(1,1) 3.55 (07) 35.11 (01) 10.41 (01) 27.27 (01) 79.99 (14) 135.37 (08) 95.51 (14) 144.52 (14) -0.03 (03) -0.10 (07) -0.07 (13) -0.28 (09)
oDistReg 3.40 (03) 49.54 (07) 11.55 (07) 39.18 (07) 74.13 (13) 143.14 (14) 91.18 (13) 132.34 (13) -0.03 (06) 0.03 (05) 0.00 (01) -0.10 (01)
oDistReg+GC 3.40 (02) 46.18 (05) 11.02 (04) 35.96 (05) 56.46 (06) 114.50 (05) 71.92 (05) 102.66 (06) -0.04 (09) 0.02 (03) -0.02 (06) -0.15 (04)
oDistReg+spGC 3.39 (01) 46.22 (06) 11.04 (05) 36.01 (06) 56.44 (05) 114.55 (06) 71.96 (06) 102.65 (05) -0.03 (06) 0.03 (05) -0.02 (05) -0.13 (03)
oMvDistReg(t, CD, OLS, ind) 3.55 (08) 60.09 (09) 13.23 (09) 48.49 (09) 71.37 (12) 139.14 (11) 88.52 (12) 126.91 (11) -0.05 (11) -0.15 (08) -0.04 (09) -0.32 (11)
oMvDistReg(t, CD, OLS) 3.62 (13) 67.75 (15) 14.95 (14) 54.82 (15) 56.16 (04) 111.61 (04) 70.22 (04) 100.64 (04) -0.03 (06) -0.03 (04) -0.01 (03) -0.18 (06)
oMvDistReg(t, CD, LASSO) 3.62 (14) 67.69 (14) 15.09 (16) 54.60 (12) 56.03 (03) 111.15 (03) 69.96 (03) 100.29 (03) -0.03 (03) -0.00 (01) -0.00 (02) -0.16 (05)
oMvDistReg(t, MCD, OLS, ind) 3.52 (04) 60.06 (08) 13.14 (08) 48.35 (08) 70.90 (10) 139.46 (12) 88.10 (10) 126.46 (10) -0.03 (03) -0.18 (11) -0.02 (06) -0.27 (08)
oMvDistReg(t, MCD, OLS) 3.61 (12) 68.64 (16) 15.00 (15) 56.01 (16) 55.39 (02) 110.49 (02) 69.26 (02) 99.20 (02) -0.06 (13) -0.15 (09) -0.05 (11) -0.26 (07)
oMvDistReg(t, MCD, LASSO) 3.62 (15) 67.62 (13) 14.91 (13) 54.65 (13) 55.32 (01) 110.23 (01) 69.20 (01) 98.83 (01) -0.02 (01) -0.01 (02) 0.01 (04) -0.12 (02)
oMvDistReg(t, LRA, OLS, ind) 3.57 (09) 60.27 (10) 13.60 (10) 48.83 (10) 71.12 (11) 140.08 (13) 88.50 (11) 126.99 (12) -0.05 (11) -0.17 (10) -0.05 (11) -0.31 (10)
oMvDistReg(t, LRA, OLS) 3.58 (10) 66.68 (12) 14.42 (12) 54.65 (14) 69.45 (09) 138.95 (10) 86.85 (09) 125.42 (09) -0.04 (10) -0.33 (13) -0.05 (11) -0.45 (13)
oMvDistReg(t, LRA, LASSO) 3.58 (11) 66.32 (11) 14.28 (11) 54.29 (11) 69.41 (08) 138.54 (09) 86.73 (08) 125.26 (08) -0.03 (06) -0.21 (12) -0.03 (08) -0.35 (12)
Table 5: Price spike and spread analysis for the CRPS and the miscoverage $\text{MC}_{0.95}$. A spike event is defined if the min/max price of a day exceeds the 5% resp. 95%-quantile of all min/max prices. A large spread event is defined as the min-to-max spread exceeding 90%-quantile of all min-to-max spreads. Numbers in the brackets give the model ranking.
Before concluding the results, we want to emphasize some limitations this study.
  • 1.
    We acknowledge that the proposed models yield a trade-off between accuracy in the marginal distributions and correct modeling of the dependence structure. This trade-off is reflected also in the disconnect between marginally dominated scoring rules (CRPS and ES) and the LS, DSS and VS and therefore, we yield somewhat ambiguous results.
  • 2.
    While the results of the spike analysis are encouraging, the parametric assumption on the distributional family might be too restrictive in some cases and could be complemented with non-parametric approaches, or with multivariate extreme value models to capture the tail behavior. This could also help to improve the forecasting performance in the marginal distributions and hence CRPS and ES.
Overall, our results highlight that neglecting the dependence structure by relying solely on marginal, univariate models yields subpar probabilistic forecasting performance. We note that for the truly multivariate approaches, using both, Copula-based combinations of univariate models and the fully multivariate distributional regression yield statistically significant performance improvements. In this regard, this paper is the first to carry out a comprehensive, multivariate probabilistic forecasting study on the day-ahead market, including also the challenging years of the COVID-19 pandemic and the energy crisis following the Russian invasion of Ukraine in the test set.
4.3 Overfitting and Regularization
Figure 14: Analysis of Overfitting. We estimate the oMvDistReg($t$,MCD,OLS) in a 8-fold cross-validation setting with 100 out-of-sample days without early stopping and monitor the in-sample and out-of-sample LS. Confidence bands are 95%-confidence intervals based on the standard deviation of the LS. The number of model coefficients increases with the amount of modeled elements. The information-criterion based early stopping suggest to include one (BIC, HQC) respectively four (AIC) off-diagonals.
Modeling all elements of the scale matrix can lead to overfitting, especially in the high dimensional setting of energy markets. To counter this, we have proposed a path-based regularization approach in Section 2.6. Figure 14 shows the results of a small experiment on the initial training set. We estimate the oMvDistReg($t$,MCD,OLS) in a 8-fold cross-validation setting with 100 out-of-sample days without early stopping. We monitor the in-sample and out-of-sample LS for the number of off-diagonals of $\bm{\mathrm{L}}$ modeled in the path-wise estimation of $\bm{\mathrm{\Omega}}=\bm{\mathrm{L}}^{\top}\bm{\mathrm{D}}\bm{\mathrm{L}}$. We see that the out-of-sample LS barely increases after modelling the first off-diagonal and starts to degrade after the 4th to 5th off-diagonal is included in the model for $\bm{\mathrm{\Omega}}$. At the same time, the number of model coefficients increases with the amount of modeled elements. The information-criterion based early stopping suggest to include one (BIC and HQC) respectively four (AIC) off-diagonals. Figure 15 shows the out-of-sample LS for the oMvDistReg($t$,MCD,OLS) model for different levels of path-based regularization, which are well-aligned with the in-sample analysis, as the LS plateaus after including the first off-diagonal. In combination with the results from the previous section, we see that our approach allows for parsimonious, yet interpretable time-varying modeling of the dependence structure by exploiting the natural ordering of the hours in the day-ahead market.
Figure 15: Results for fitting the path-based regularization for the oMvDistReg($t$,MCD,OLS) model.
4.4 Computation Time
Model Initial Fit Avg. Update Total Time Est. Speedup
LARX+N(0, $\sigma$) 1.23 0.06 103.80 $\times$ 21
LARX+N(0, $\Sigma$) 1.23 0.06 103.80 $\times$ 21
LARX+Adaptive CP 1.23 0.06 103.80 $\times$ 21
LARX+GARCH(1,1) 1.44 0.27 501.31 $\times$ 5
oDistReg 55.31 0.30 598.78 $\times$ 169
oDistReg+GC 55.32 0.30 599.61 $\times$ 168
oDistReg+spGC 55.46 0.45 870.45 $\times$ 116
oMvDistReg(t, CD, OLS, ind) 38.26 0.07 173.88 $\times$ 402
oMvDistReg(t, CD, OLS) 123.11 0.18 460.38 $\times$ 489
oMvDistReg(t, CD, LASSO) 460.46 2.34 4736.02 $\times$ 178
oMvDistReg(t, MCD, OLS, ind) 41.72 0.06 147.07 $\times$ 519
oMvDistReg(t, MCD, OLS) 132.20 0.15 403.89 $\times$ 599
oMvDistReg(t, MCD, LASSO) 1484.75 1.79 4754.88 $\times$ 571
oMvDistReg(t, LRA, OLS, ind) 21.82 0.03 83.58 $\times$ 477
oMvDistReg(t, LRA, OLS) 289.85 0.77 1702.47 $\times$ 311
oMvDistReg(t, LRA, LASSO) 5654.00 3.73 12473.45 $\times$ 829